OpenStudy (anonymous):

To decorate a cake, you make a shade of green icing using 3 drops of yellow food coloring for every 4 drops of blue. If you use 12 drops of yellow food coloring, how many drops of blue food coloring would you need?

OpenStudy (anonymous):

\[3x \div60=20\]

OpenStudy (anonymous):

So 20 drops of blue food coloring

OpenStudy (anonymous):

yellow 3 12 ------ = ----- = ----- blue 4 x cross multiply (3)(x) = (12)(4) 3x = 48 x = 48/3 x = 16 so, 3/4 = 12/16 This is a proportion, and basically a proportion is nothing but equivalent fractions. As you can see...12/16 reduces to 3/4.

OpenStudy (anonymous):

that would be 16 drops of blue food coloring needed for 12 drops of yellow.

OpenStudy (goformit100):

A Warm Welcome To Open Study